Given the following parameters of these MOSFETS gm = 500 x 10-6 A/V, ro = ∞

A. Calculate the loop Gain - Aβ

B. Determine the oscillation frequency ω o

C. The Rd value required for oscillation

Rd 20 R

Pls explain each step in details and clear results
Answer #1

& dila RE HhMh 2c R - Vor - gm Rd [ - Bil]Now B= T ac e. TEM FR Z2 = zie Z1 + Z2 s Y YN 리 Zi= ell Zi = 21 - verb IASIR It ZSIR and 72 = R+ asc. - - 254. R NOW Itsch vocompanity imaginary funn in 6 (9m Rd) WRC = 5wRc. gm Rd=5 5 Rd=- 500 x 10-6 |Rd=10kn (B) compart, Real tama in co. 1-2 w2e2c2